OpenStudy (anonymous):

i need help solving this 5x -7y=9 4x + 3y =33

OpenStudy (jim766):

you need to get rid of one of the variables... x or y in this case it doesnt matter which one, what do you want to get rid of?

OpenStudy (anonymous):

the y

OpenStudy (jim766):

we need to make it so when we add the two equations together, the y variable will be eliminated.

OpenStudy (jim766):

we can multiply a number times an equation....as long as we multiply that number times everything in the equation.

OpenStudy (jim766):

In this problem, the best way to do that is to 1. multiply the top equation by the coefficient of the bottom y (3) 2. multiply the bottom equation by the coefficent of the top y (7) it would look like this: (3) 5x - 7y = 9 (7) 4x + 3y = 33 15x - 21y = 27 notice how now the coeff are the same number 28x + 21 y = 231 but with different signs...when we add them we get 0y which is 0 43x = 258 now divide both sides by 43 to find x. once you what x equals, substitute that into either of the original equations and solve for y try to finish it on your own